2-(N-morpholino)ethane sulfonic acid (commonly referred to as MES) is widely used as a pH buffer for biochemical experiments. It has a pKa of 6.15 and is usually purchased in the acidic form. Suppose that you were to dissolve 0.2 moles of MES in 1 L of water and then add to this 150 mL of 1 M NaOH. Calculate the pH of the resulting solution.
Answer #1

Answer.

The above reaction can be written as

MES + NaOH ---> MES (salt) + H2O

Given that, pKa = 6.15

Molarity of MES = volume*molarity = 0.2*1 = 0.2 moles

Moles of NaOH = volume*molarity = 0.15*1 = 0.15 moles

Therefore,

MES + NaOH ---> MES (salt) + H2O

t=0 0.2 0.15 - -

t = teq 0.05 - 0.15

Using Henderson-Hasselbalch Equation For Acid Buffer,

pH = pKa + log(salt/acid)

pH = 6.15 + log(0.15/0.05)

pH = 6.15 + log(3) = 6.627

Therefore, pH of solution is 6.627.